# A comprehensive theory for relativistic polaritonic chemistry: a way toward a four component ab initio treatment of molecular systems

### v1

@@ II.2.4 Treatment of the negative energy states
Physically, the negative energy solutions correspond to positronic states. In relativistic quantum chemistry, the negative energy states are usually filled with electrons, therefore, no electron can fall in these states. However, if one electron is removed from the negative energy states, it would leave a hole corresponding to the creation of a positively charged particle, the positron. This reinterpretation becomes clearer if we split the sum over all state into a sum over the positive energy states and another over the negative energy ones.In this picture, the negative energy electron creation (annihilation) operators are reinterpreted as positive energy positron annihilation (creation) operators (see Ref.s [60, 61, 62, 56]). In this framework, the field operator can be decomposed in this way:
@@
and and the modified one and two-electrons integrals have been defined in Eq.47. Barred indices correspond to positronic indices.
@@ II.3 The Hartree-Fock approximation
Since it treats explicitly all the interactions between relativistic electrons/positrons and the photons of the cavity field, Hamiltonian 51 represents the perfect starting point for the development of ab initio theories that can be used to simulated molecular systems containing heavy atoms in optical cavities. In this section, we will use Hamiltonian 51 to develop the first relativistic 4-components Hartree-Fock (HF) approach for polaritonic chemistry. In quantum chemistry, HF represents the simplest approximation respecting the right symmetry of all the involved particles. HF also provides access to physically meaningful sets of atomic/molecular orbitals that can be subsequently used to develop more accurate correlated theories. HF is based on the independent particle approximation, meaning that the involved quantum species do not directly interact between each other but travels in space under the action of the average potential of the other particles (mean field approximation). For the purely electronic case, this implies that the wave function is described by a single Slater determinant of atomic/molecular orbitals or spinors in the relativistic case:
This ansatz ensures that the anti-symmetry of the wave function for a particle exchange will be respected. In this case we deal with three kinds of particles: electrons, positrons and photons, each one respecting its own statistics (fermionic for electrons and positrons and bosonic for photons). The independent particles condition, in this situation, can be enforced defining the wave function ansatz as the product of distinct wave functions for every species:
where and are single determinants for electronic and positronic spinors respectively, is the vacuum state associated to the photonic mode . A similar ansatz has been already applied in the non-relativistic version of the QED-HF approach, and its implications are discussed in detail in Ref. [24]. Ansatz 55 can be rewritten, for practical reasons, in terms of second quantized electronic and positronic operators as:
@@
Notice that at the HF level, all terms that do not conserve the number of particles will have a zero expectation value. This is the case for the so-called bilinear term in Eq.s 46 and 51 which involves only one annihilation (creation) operator and therefore changes the number of photon. Consequently, this term gives a zero contribution to the ground state energy.
@@ II.3.1 The Fock-operator
where are either positron or electron indices. Hence the Fock operator in block-matrix form is the following:
where indices () correspond to occupied(unoccupied) electronic spinors. Barred indices () refer instead to positronics orbitals. The Fock operator in Eq. 68 can be applied in a Roothan-Hall like procedure [63]:
@@
In standard relativistic quantum chemistry, the positrons are usually removed from the treatment by application of the no-pair approximation [57, 45]. Consequently, the spinor optimization procedure is in fact an excited state problem where the ground state energy corresponds to the first positive energy state. This is referred to as a minimax problem [57, 45]. In our approach, equivalent results can be obtained when no explicit positron is included in the wave function ansatz:
@@ II.3.2 Kinetic Balance in presence of the field
In relativistic quantum chemistry, without including positrons, the spinors can be expressed as two-components objects:
where and are called respectively ”Large” and ”Small” components. In practice, the spinor solutions are expanded on a basis-set: where and is a 2-spinor. In order to avoid the so-called varitional collapse of the solution, a constraint on the large and small components of the basis set is applied [53, 45]. In the presence of a vector potential, such constraint has the following form:
where is the vector of the three Pauli matrices and is the generalized momentum. Such prescription is called the magnetic balance condition [64, 65, 66, 67]. In standard relativistic quantum chemistry, the momentum is simply . We refer to this condition as the kinetic balance [53, 45]. In our case, since the generalized momentum is a priori depending on the field, the magnetic balance needs to be in principle satisfied. This could require to have a field-dependent small-component. However, in the present context, we have shown that in the dipole-approximation, applying the length gauge transformation, the momentum of the theory is therefore transformed in the following way (see App.A):
@@
can be applied in this case. It is important to stress that this is only possible under the dipole-approximation, otherwise, the length gauge transformation would not provide such a simple shape for the Hamiltonian and the associated momentum.
@@ II.4 Excited states properties
In polaritonic chemistry, the signature property emerging from the strong coupling condition is the Rabi-splitting. It represents the energy separations between the polaritons formed by the mixing between matter and field states (see Fig. 2).
Calculating Rabi splittings, which could directly be compared with experimental data, requires access to the excited states of the coupled light matter system. At the HF level they can be simulated recurring to linear response theory. Recently, Castagnola et al. [68] presented an HF linear response theory for non-relativistic polaritonic systems. In this section, we present an extension of this approach to 4-components Dirac HF.Using linear response theory, excitations energies () can be obtained by solving the well known Casida equation [69]:
The polaritonic espressions of the and matrices as derived in Ref. [68] assumes the form:
@@
A similar approach, has been recently proposed by Konecny et al. [30] to calculate excitation energies at the QED Dirac Kohn Sham level of theory.
@@ III Results
All the results presented in this section have been obtained from the development version of the PySCF software package [70, 71]. In particular, the Pol-DHF approach has been implemented by applying modifications to the integrals used in standard DHF [72]. So far, to simplify the interpretation of the results, we neglected the positrons in the treatment. A detailed discussion of the effects induced by the explicit inclusion of the positronic degrees of freedom will the topic of a future follow-up paper. All calculations have been performed using the all-electron x2c-SVP basis-set [73, 74, 75, 76] for the large component. The small component has been obtained via the kinetic balance prescription. The coupling value has been set to 0.05 a.u..
@@ III.1 Ground state properties
In this section, we analyze the field induced effects on the ground state properties of three metal hydrides (CuH, AgH and AuH). These complexes contain metals belonging to different periods of the 11th group of the periodic table. Going down the groups, the velocity of the electrons (in particular of the inner ones) increases, approaching finite fractions of the speed of light in the gold case. In fact, gold complexes are well known to have peculiar electronic properties due to the significant relativistic effects [77, 78, 79, 80]. In this paper, we will analyze how the polaritonic effects compete with the relativistic ones generating modification of the molecular electronic structure. In Table 1 we compare the total energies calculated at the DHF level of theory with those evaluated including the field (Pol-DHF).
@@
The field induced effects, despite small compared to the total energies, represent still a sizable (some tens of eV) variation on the energy of the system. In particular, it is interesting noticing that the field induced energy variation is quite similar for copper and gold (Cu - 0.15 eV, Au - 0.20 eV) while it is slightly bigger (in absolute value) for silver (Ag - 0.24 eV). This trend could be actually explained if we realize that under the QED-HF approximation, the only field term contributing to the total energy of the system is the dipole self-energy contribution . As highlighted in Eq. 47 the one-electron term coming from the dipole self-energy directly depends on the molecular quadrupole, and it is well known from the literature (see Ref. [81]) that this as many other electronic properties (atomic radius, etc.) [82] show a very similar trend if we move down the group.
@@
In Figure 3 we compare the orbital energies for the valence orbitals ( and ) of the three complexes evaluated at the Pol-DHF level with those calculated at the DHF and scalar 2-components X2C level of theory [83, 70, 71]. As expected, the biggest variations in the orbital energies can be observed comparing the scalar X2C results with the DHF ones. in this case, the usual destabilization of the orbitals and consequent stabilization of the MOs can be observed. The reduction of the gap between these occupied MOs decreases with the increasing relativistic character of the metal (smaller for Au than Cu). Much smaller, even if sizable, is the effect produced by the field.
@@
The field induce effects () on the energy gap, barely appreciable in Figure 3, are reported in Table 2. As already observed for the total energy also in the field induced variation of the energy gap is larger for AgH compared to the CuH and, in particular, to AuH. It is crucial, to point out that also in the gold case (smallest effect) the cavity field is able to induce a 10 meV variation on the orbital gap. These values, represent a reasonable fraction ( or higher) of a kcal/mol, relevant to observe variations of the chemical properties. Even more interesting is observing that the relative change induced by the field () is larger for AuH (about 1% of the total gap) compared to the other systems (0.7% for AgH and 0.6% for CuH). Notice that these field induced variations of the energy gap bring (for all the systems) to a reduction of the gap due to a very small stabilization of the MOs accompanied by a larger destabilization of the orbitals.
@@ III.1.1 Gaunt, Breit and polaritonic contributions to ground state energies
As mentioned in section II.2.1 the radiative QED corrections and in particular the retardation effects can compete or sum up to the polaritonic contributions in a relativistic molecular system. Here we perform a detailed comparison between the energy variations introduced by quantum field, with the Gaunt and Breit terms contributing to the full retarded potential.
@@
In Table 3 the Gaunt/Breit corrections to the DHF and Pol-DHF energies are presented. These effects are quite sizable, in particular for AuH that is the system showing the largest relativistic effects. In this case the trend is monotonic and both the Gaunt and Breit effect increase going down the group of the periodic table. It is important to notice that if the absolute energy variation is taken into account, these effects are at least one order of magnitude larger than the effects generated by the quantum field. Emblematic is the comparison between the Breit correction in CuH (the less relativistic system) and the corresponding polaritonic energy correction, shown in Table 1. These observations clearly indicate that, for a molecular system coupled to a quantum field, attention needs to be paid in omitting the retardation effects, a generic pratice in standard electronic relativistic quantum chemistry simultations.More striking results can be obtained if we perform the same analysis on the orbital energies. In Figure 4 the energy contributions (in logaritmic scale) due to the Gaunt, Breit and polaritonic terms on the energies of the occupied MOs are presented for the three systems.
@@
The most evident aspect is the very wide variation range of the energy contributions for different molecular orbitals. In general, the effects due to the Full-Breit term are much more sizable for the core orbitals, while they monotonically decrease moving toward the valence ones. This trend is clearly expected, since the inner orbitals host the fastest electrons that are more affected by the relativistic effects. This behavior is obviously more evident for AuH than for CuH. The polaritonic contribution shows instead much smaller variations, and the effect slightly increases moving toward the valence. In the core region, the retardation effects are always larger, while in the valence region the polaritonic effects are dominant. It is interesting highlighting that for the intermediate orbitals all the effects are quite comparable, confirming that the inclusion of retardation effects could produce crucial effects in molecular systems coupled to photons. There are anyway cases where, neglecting such effects, as usually done in relativistic quantum chemistry, is still reasonable. For instance, if we are interested in excited states properties (i.e. Rabi Splitting, etc.) involving only valence electrons, a phenomenon recently analyzed by Konecny et al. in Ref. [30], then omitting retardation effects in the treatment is clearly acceptable. However, if excitations from inner orbitals (i.e. core excitations, etc.) or magnetic properties of the system need to be analyzed the inclusion of the retardation effects will become a crucial point in the treatment.
It is important to point out that the implications of the effects presented in this section are, in some sense, minimized by the absence of electron-electron and electron-photon correlation in the treatment. For instance, the HF approximation removes the frequency dependence of the ground state energy from the frequency of the field. This dependence can be recovered only by including correlation into the model. We expect, that the inclusion of frequency dependent terms, will be crucial in particular to describe resonant processes. The inclusion of electron-electron and electron-photon correlation will be the main topic of a future follow-up paper.
@@ III.2 Excited state properties
In the section III.1, we focused on the field induced effects on the ground state properties of three metal hydrides of the 11th group of the periodic table. In this section, we analyze in details the effects generated by the photons on the optical properties of the AuH complex [84, 85]. Similar analysis for CuH and AgH is presented in Appendix B. In this system, because of strong relativistic effects, we expect to observe spectra very different compared to those that can be simulated without taking relativity into account. This is evident if we look at Figure 5.
As it can be seen, at the HF level, the triplet state () is found to be higher in energy than the singlets states, in disagreement with the experimental data from Ref.s [84, 85]. The inclusion of scalar relativistic correction partially resolve this issue, but it underestimates the energy by about 0.45 eV compared to DHF calculations. DHF does not only reproduce the right ordering of the states, but it also describes the breaking of the degeneracy to form (from the triplets) the and states. For an improved readability of the plot Figure 5 the Pol-DHF data have been calculated off-resonance. This choice highlights the energy shift due to the dipole self energy. In resonance conditions also the Rabi splitting (discussed later in Figure 6) would be visible.
Comparing this data with the bare electronic DHF ones, we see that the field (in off resonance conditions) induces a stabilization of the and states but without inducing any change in the ordering of the states.
@@
This spin-orbit coupling induced singlet-triplet rabi-splitting was already reported for a different system by Koneckny et al. in Ref. [30]. These observations clearly prove how the electromagnetic field can be used to manipulate and control inter-system crossing processes and consequently the phosphorescence of complexes containing heavy atoms. In App. B a similar discussion is reported also for CuH and AgH.

### v2

@@ II.2.4 Treatment of the negative energy states
Physically, the negative energy solutions correspond to positronic states [67]. This reinterpretation becomes clearer if we split the sum over all state into a sum over the positive energy states and another over the negative energy ones. In this picture, the negative energy electron creation (annihilation) operators are reinterpreted as positive energy positron annihilation (creation) operators (see Ref.s [67, 68, 69, 61]). In this framework, the field operator can be decomposed in this way:
@@
and and the modified one and two-electrons integrals have been defined in Eq.47. Barred indices correspond to positronic indices. In this article we do not address the role of the fermionic vacuum that is at the heart of many discussions [68, 69, 61]. Indeed, the choice of the Hamiltonian and the possible reinterpretation of negative energy states has an influence on the nature of the fermionic vacuum which in turn has some consequence on the expectation value of the Hamiltonian. Fortunately, at the Hartree-Fock level there is no real dependence on the vacuum nature (however, it becomes important when developing correlated methods).
@@ II.3 The Hartree-Fock approximation
Since it treats explicitly all the interactions between relativistic electrons/positrons and the photons of the cavity field, Hamiltonian 51 represents the perfect starting point for the development of ab initio theories suitable to simulate molecular systems containing heavy atoms in optical cavities. In this section, we will use Hamiltonian 51 to develop the first relativistic 4-components Hartree-Fock (HF) approach for polaritonic chemistry. In quantum chemistry, HF represents the simplest approximation respecting the right symmetry of all the involved particles. HF also provides access to physically meaningful sets of atomic/molecular orbitals that can be subsequently used to develop more accurate correlated theories. HF is based on the independent particle approximation, meaning that the involved quantum species do not directly interact between each other but travels in space under the action of the average potential of the other particles (mean field approximation). For the purely electronic case, this implies that the wave function is described by a single Slater determinant of atomic/molecular orbitals or spinors in the relativistic case:
This ansatz ensures the anti-symmetry of the wave function under particle exchange. In this case, three kinds of particles are involved: electrons, positrons and photons, each one respecting its own statistics (fermionic for electrons and positrons and bosonic for photons). The independent particles condition, in this situation, can be enforced defining the wave function ansatz as the product of distinct wave functions for every species:
where and are single determinants for electronic and positronic spinors respectively, is the vacuum state associated to the photonic mode and is a shorthand notation for . A similar ansatz has been already applied in the non-relativistic version of the QED-HF approach, and its implications are discussed in detail in Ref. [24]. Ansatz 55 can be rewritten, for practical reasons, in terms of second quantized electronic and positronic operators as:
@@
Notice that at the HF level, all terms that do not conserve the number of particles will have a zero expectation value. This is the case for the so-called bilinear term in Eq.s II.2.3, 46 and 51 which involves only one annihilation (creation) operator and therefore changes the number of photon. Consequently, this term gives a zero contribution to the ground state energy.
@@ II.3.1 The Fock-operator
where are either positron or electron indices. The Fock operator in Eq. 68 can be applied in a Roothan-Hall like procedure [70]:
@@
In standard relativistic quantum chemistry, the positrons are usually removed from the treatment by application of the no-pair approximation [62, 53]. Consequently, the spinor optimization procedure is in fact a minimax problem [62, 53]. In our approach, equivalent results can be obtained when no explicit positron is included in the wave function ansatz:
@@ II.3.2 Kinetic Balance in presence of the field
In relativistic quantum chemistry, the spinors can be expressed as two-components objects:
where and are called respectively ”Large” and ”Small” components. In practice, the spinor solutions are expanded on a basis-set: where and is a 2-spinor. In order to avoid the so-called variational collapse of the solution, a constraint on the large and small components of the basis set is applied [58, 53]. In the presence of a vector potential, such constraint has the following form:
where is the vector of the three Pauli matrices and is the generalized momentum. Such prescription is called the magnetic balance condition [71, 72, 73, 74]. In standard relativistic quantum chemistry, the momentum is simply . We refer to this condition as the kinetic balance [58, 53]. In our case, since the generalized momentum is a priori depending on the field, the magnetic balance needs, in principle, to be satisfied. This could require to have a field-dependent small-component basis-set. However, in the present context, we have shown that in the dipole-approximation, applying the length gauge transformation, the momentum of the theory is thus transformed in the following way (see App.A):
@@
can be applied in this case. It is important to stress that this is only possible under the dipole-approximation, otherwise, the length gauge transformation would not provide such a simple expression for the Hamiltonian and the associated momentum.
@@ II.4 Excited states properties
In polaritonic chemistry, the signature property emerging from the strong coupling condition is the Rabi-splitting. It represents the energy separation between the polaritons formed by the mixing between matter and field states (see Fig. 2).
Calculating Rabi-splittings, which could directly be compared with experimental data, requires access to the excited states of the coupled light matter system. At the HF level they can be simulated recurring to linear response theory. Recently, Castagnola et al. [75] presented an HF linear response theory for non-relativistic polaritonic systems. In this section, we present an extension of this approach to 4-components Dirac HF. Using linear response theory, excitations energies () can be obtained by solving the well known Casida equation [76]:
The polaritonic expressions of the and matrices as derived in Ref. [75] assumes the form:
@@
A similar approach, has been recently proposed by Konecny et al. [35] to calculate excitation energies at the QED Dirac Kohn Sham level of theory based on previous work presented in the following references [77, 78].
@@ III Results
All the results presented in this section have been obtained from the development version of the PySCF software package [79, 80]. In particular, the Pol-DHF approach has been implemented by applying modifications to the integrals used in standard DHF [81]. So far, to simplify the interpretation of the results, we neglected the positrons in the treatment. A detailed discussion of the effects induced by the explicit inclusion of the positronic degrees of freedom will the topic of a future follow-up paper. All calculations have been performed using the all-electron x2c-SVP basis-set [82, 83, 84, 85] for the large component. The small component has been obtained via the kinetic balance prescription. The coupling value has been set to 0.05 a.u.. The datasets generated and analyzed during the current study can be reproduced using the PySCF development code, the inputs files and the geometries provided at the Zenodo link [86].
@@ III.1 Ground state properties
In this section, we analyze the field induced effects on the ground state properties of three metal hydrides (CuH, AgH and AuH). These complexes contain metals belonging to different periods of the 11th group of the periodic table. Going down the groups, the velocity of the electrons (in particular of the inner ones) increases, approaching finite fractions of the speed of light in the gold case. In fact, gold complexes are well known to have peculiar electronic properties due to the significant relativistic effects [87, 88, 89, 90]. In this paper, we will analyze how the polaritonic effects compete with the relativistic ones, generating modification of the molecular electronic structure. In Table 1 we compare the total energies calculated at the DHF level of theory with those evaluated including the field (Pol-DHF).
@@
The field induced effects, despite small compared to the total energies, still represent a sizable (some tens of eV) variation on the energy of the system. In particular, it is interesting noticing that the field induced energy variation is quite similar for copper and gold (Cu - 0.15 eV, Au - 0.20 eV) while it is slightly bigger (in absolute value) for silver (Ag - 0.24 eV). This trend can actually be explained by the fact that under the QED-HF approximation, the only field term contributing to the total energy of the system is the dipole self-energy contribution . As highlighted in Eq. 47 the one-electron term coming from the dipole self-energy directly depends on the molecular quadrupole, and it is well known from the literature (see Ref. [91]) that this as many other electronic properties (atomic radius, etc.) [92] show a very similar trend if we move down the group.
@@
In Figure 3, we compare the orbital energies for the valence orbitals ( and ) of the three complexes evaluated at the Pol-DHF level with those calculated at the DHF and scalar 2-components X2C level of theory [93, 79, 80]. As expected, the biggest variations in the orbital energies can be observed comparing the scalar X2C results with the DHF ones. In this case, the usual destabilization of the orbitals and consequent stabilization of the MOs can be observed. The reduction of the gap between these occupied MOs decreases with the increasing relativistic character of the metal (smaller for Au than Cu). The effect of the field on the electronic structure, though sizeable, is much smaller.
@@
The field induced effects () on the energy gap, barely appreciable in Figure 3, are reported in Table 2. As already observed for the total energy, the largest field induced variation of the energy gap is observed for AgH. It is crucial, to point out that even in the gold case, where the effect is smaller, the cavity field is able to induce a 10 meV variation on the orbital gap. These values, represent a reasonable fraction ( or higher) of a kcal/mol, relevant to observe variations of the chemical properties. Even more interesting is observing that the relative change induced by the field () is larger for AuH (about 1% of the total gap) compared to the other systems (0.7% for AgH and 0.6% for CuH). Notice that these field induced variations of the energy gap lead (for all systems) to a reduction of the gap due to a very small stabilization of the MOs accompanied by a larger destabilization of the orbitals.
@@ III.1.1 Gaunt, Breit and polaritonic contributions to ground state energies
As mentioned in section II.2.1, the retardation effects can compete or sum up to the polaritonic contributions in a relativistic molecular system. Here we perform a detailed comparison between the energy variations introduced by quantum field, with the Gaunt and Breit terms contributing to the full retarded potential.
@@
In Table 3, the Gaunt/Breit corrections to the DHF and Pol-DHF energies are presented. These effects are quite sizable, in particular for AuH which is the system exhibiting the largest relativistic effects. In this case, the trend is monotonic and both the Gaunt and Breit effect increase going down the group of the periodic table. It is important to notice that if the absolute energy variation is taken into account, these effects are at least one order of magnitude larger than the effects generated by the quantum field. The comparison between the Breit correction in CuH (the less relativistic system) and the corresponding polaritonic energy correction, shown in Table 1, demonstrates that quite unequivocally. These observations clearly indicate that, for a molecular system coupled to a quantum field, attention needs to be paid in omitting the retardation effects, a generic practice in standard electronic relativistic quantum chemistry simulations. More striking results can be obtained if we perform the same analysis on the orbital energies. In Figure 4, the absolute value of the energy contributions (in logarithmic scale) due to the Gaunt, Breit and polaritonic terms on the energies of the occupied MOs are presented for the three systems.
@@
The most evident aspect is the very wide variation range of the energy contributions for different molecular orbitals. In general, the effects due to the full-Breit term are much more sizable for the core orbitals, while they monotonically decrease moving toward the valence ones. This trend is clearly expected and has been documented in Ref. [63]. This behavior is obviously more evident for AuH than for CuH. The polaritonic contribution shows instead much smaller variations, and the effect slightly increases moving toward the valence. In the core region, the retardation effects are always larger, while in the valence region the polaritonic effects are dominant. Interestingly, for the intermediate orbitals, all the effects are quite comparable, confirming that the inclusion of retardation effects could induce crucial effects in molecular systems coupled to photons. Still, there are cases where neglecting such effects, as usually done in relativistic quantum chemistry, is still reasonable. For instance, if we are interested in excited states properties (i.e. Rabi Splitting, etc.) involving only valence electrons, a phenomenon recently analyzed by Konecny et al. in Ref. [35], then omitting retardation effects in the treatment is clearly meaningful. For this reason, in the next section excited states will be investigated without including the full-Breit correction. However, if excitations from inner orbitals (i.e. core excitations, etc.) or magnetic properties of the system need to be analyzed, the inclusion of the retardation effects will become necessary. Bear in mind that the implications of the effects presented in this section are, in some sense, minimized by the absence of electron-electron and electron-photon correlation in the treatment. For instance, the HF approximation removes the frequency dependence of the ground state energy from the frequency of the field. This dependence can be recovered only by including correlation into the model [27]. We expect, that the inclusion of frequency dependent terms, will be crucial in particular to describe resonant processes. The inclusion of electron-electron and electron-photon correlation will be the main topic of a future follow-up paper.
@@ III.2 Excited state properties
In the section III.1, we focused on the field induced effects on the ground state properties of three metal hydrides of the 11th group of the periodic table. In this section, we analyze in details the effects generated by the photons on the optical properties of the AuH complex [94, 95]. Similar analysis for CuH and AgH is presented in Appendix B. In this system, because of strong relativistic effects, we expect to observe spectra very different compared to those simulated without taking relativity into account. This is obvious looking at Figure 5.
As it can be seen, at the HF level, the triplet state () is found to be higher in energy than the singlets states, in disagreement with the experimental data from Ref.s [94, 95]. The inclusion of scalar relativistic correction partially resolve this issue, but it underestimates the energy by about 0.45 eV compared to DHF calculations. DHF does not only reproduce the right ordering of the states, but it also describes the breaking of the degeneracy to form (from the triplets) the and states. For an improved readability of the plot Figure 5 the Pol-DHF data have been calculated off-resonance. This choice highlights the energy shift due to the dipole self energy. In resonance conditions, the Rabi splitting (discussed later in Figure 6) would also be visible.
Comparing this data with the bare electronic DHF ones, we see that the field (in off resonance conditions) induces a stabilization of the and states without inducing any change in the ordering of the states.
@@
This spin-orbit coupling induced singlet-triplet Rabi-splitting was already reported for a different system by Konecny et al. in Ref. [35]. These observations clearly demonstrate how the electromagnetic field can be used to manipulate and control inter-system crossing processes and consequently the phosphorescence of complexes containing heavy atoms. In App. B a similar discussion is also reported for CuH and AgH.